Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 102     Entry time: 16 Mar 2020, 09:36
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes Dipatch Meeting March 16th, 2020  
FXE:
- Beamline is in safe state;
- Minor Jungfrau calibration issues over last week. In touch with DET and EEE;

SPB:
- Good exercise: experiment run with only 2 users. Successful beamtime with data taken: to be analysed by users;

MID:
- Successful beamtime with good beam conditions and data has been taken;
- Beamline is in safe state and most of staff is in telecommuting;

EEE:
- Start collecting a list of broad activities to take place in case of a long shutdown;
- No major issues to be fixed;

CTRL:
- Few cameras and detectors hiccups – e.g. DAQ re-start bring system back;
- Most of people will run from home office / telecommuting;
- Tests together with EEE/FXE still to be performed this week - to be scheduled;

DET:
- DET is placing a plan to warmed up detectors and be ready for shutdown;
- AGIPD, DSSC are already in safe mode;
ELOG V3.1.4-7c3fd00